# Unveiling the Role of Learning Rate Schedules via Functional Scaling Laws

### v1

@@ Abstract
Scaling laws have played a cornerstone role in guiding the training of large language models (LLMs). However, most existing works on scaling laws primarily focus on the final-step loss, overlooking the loss dynamics during the training process and, crucially, the impact of learning rate schedule (LRS). In this paper, we aim to bridge this gap by studying a teacher–student kernel regression setup trained via online stochastic gradient descent (SGD). Leveraging a novel intrinsic time viewpoint and stochastic differential equation (SDE) modeling of SGD, we introduce the Functional Scaling Law (FSL), which characterizes the evolution of population risk during the training process for general LRSs. Remarkably, the impact of the LRSs is captured through an explicit convolution-type functional term, making their effects fully tractable. To illustrate the utility of FSL, we analyze three widely used LRSs – constant, exponential decay, and warmup-stable-decay (WSD) – under both data-limited and compute-limited regimes. We provide theoretical justification for widely adopted empirical practices in LLMs pre-training such as (i) higher-capacity models are more data- and compute-efficient; (ii) learning rate decay can improve training efficiency; (iii) WSD-like schedules can outperform direct-decay schedules. Lastly, we explore the practical relevance of FSL as a surrogate model for fitting, predicting and optimizing the loss curves in LLM pre-training, with experiments conducted across model sizes ranging from 0.1B to 1B parameters. We hope our FSL framework can deepen the understanding of LLM pre-training dynamics and provide insights for improving large-scale model training.
@@ 1 Introduction
It has been empirically observed that the training of large-scale deep models follows certain scaling laws (hestness2017deep, ; kaplan2020scaling, ; hoffmann2022training, ): as model and dataset sizes increase, the performance improves predictably in a power-law fashion, i.e., the population loss $L$ satisfies:
$L=L_{0}+C_{M}M^{-\alpha_{M}}+C_{D}D^{-\alpha_{D}},$
where $M$ and $D$ represent the model and dataset sizes respectively, $\alpha_{M}$ and $\alpha_{D}$ are the corresponding scaling exponents, $L_{0}$, $C_{M}$ and $C_{D}$ are constants. In practice, these parameters can be estimated from smaller-scale pilot experiments and then extrapolated to accurately predict performance at larger scales touvron2023llama ; liu2024deepseek ; li2025predictablescalei .
@@
Table 1: LRS impacts the scaling efficiency in the teacher-student kernel regression. Easy-learning regime: $s\!\geqslant\!1-1/\beta$; Hard-learning regime: $s\!<\!1-1/\beta$, where $s$ gauges task difficulty and $\beta$ reflects model capacity.
@@
Exponential-decay $(D/\log D)^{-\frac{s\beta}{1+s\beta}}$ $(D/\log D)^{-s}$ $(C/\log C)^{-\frac{s\beta}{2+s\beta}}$ $(C/\log C)^{-\frac{s\beta}{1+{\beta}}}$
@@
Scaling laws have since become foundational principles in guiding the training of modern LLMs (henighan2020scaling, ; zhai2022scaling, ; kadra2023power, ; aghajanyan2023scaling, ; bi2024deepseek, ; shuai2024scaling, ; kumar2024scaling, ). Practitioners heavily rely on these laws to optimally allocate computational resources between data volume and model size prior to training hoffmann2022training , as well as to tune optimization hyper-parameters such as learning rates and batch sizes liu2024deepseek ; li2025predictablescalei .
Given the cornerstone role of scaling laws in developing modern LLMs, explaining how these laws emerge – especially how the scaling exponents $\alpha_{M}$ and $\alpha_{D}$ depend on factors such as task difficulty and model capacity – has become a central topic of modern learning theory. Recently, a growing body of work has tackled this problem from diverse theoretical perspectives (sharma2020neural, ; hutter2021learning, ; maloney2022solvable, ; wei2022more, ; jain2024scaling, ; michaud2024quantization, ; nam2024exactly, ; atanasov2024scaling, ; dohmatob2024tale, ; bahri2024explaining, ; bordelon2024dynamical, ; lin2024scaling, ; paquette2024fourplus, ; bordelon2024feature, ; zhang2024does, ). It is worth noting that unlike classical worst‑case generalization‑bound analyses shalev2014understanding , studying scaling laws demands explicitly tracking the learning trajectory for each instance. For this reason, nearly all recent theoretical studies adopt a teacher–student kernel regression setup due to its analytical tractability – a setting we also follow in this work.
A key limitation of existing theoretical studies is their focus on the scaling behavior of the final-step risk, while overlooking the loss dynamics during training process and, crucially, the role of the learning rate schedule (LRS). Yet, it is well established that LRS significantly affects training efficiency in deep learning (bengio2012practical, ; mccandlish2018empirical, ; bergsma2025straight, ; hagele2024scaling, ). As a result, different LRSs can induce distinct scaling behaviors, with the scaling exponent itself potentially depending on the schedule. Developing a theoretical framework that captures the influence of LRS on scaling laws is not only of academic interest, but also practically relevant for designing effective training strategies.
1.1 Our Contribution
In this paper, we provide the first systematic study of the interaction between LRS and scaling laws. Working on the setup of teacher–student kernel regression – widely used in theoretical studies of scaling laws (maloney2022solvable, ; bordelon2024dynamical, ; bahri2024explaining, ; paquette2024fourplus, ; lin2024scaling, ; bordelon2024feature, ), we introduce a novel Functional Scaling Law (FSL) that characterizes the risk dynamics under general LRSs. The impact of the schedules is captured through an explicit convolution-type functional term, making their effects fully transparent (Theorem 4.1).
To demonstrate the applicability and insights offered by FSL, we analyze three widely used LRSs, investigating their influence on scaling laws in both the traditional data-limited and modern compute-limited regimes. Our key findings are summarized as follows:
  • Constant LRS. We recover prior results (bordelon2024dynamical, ; bordelon2024feature, ; paquette2024fourplus, ) and extend them by incorporating the effect of label noise, which models the irreducible risk presented in real-world LLM pre-training. Notably, compute‑optimal training demands a learning rate that decreases with compute budget and higher‑capacity models are more compute‑efficient, matching empirical observations (bi2024deepseek, ; shuai2024scaling, ).
  • Exponential-decay LRS. We recover and refine the results of (lin2024scaling, ), now incorporating the effect of batch size. In particular, we show that exponential decay significantly boosts the scaling exponent compared to the constant LRS, providing a theoretical explanation for the cruciality of learning rate decay; see Table 1.
  • Warm-stable-decay (WSD) LRS. We also analyze a WSD-like LRS that maintains a constant learning rate for a prolonged stable phase before applying exponential decay near the end of training. We reveal that such schedules can substantially outperform direct exponential decay in terms of optimal scaling efficiency (see Table 1), offering theoretical justification for the widespread empirical success of WSD strategies (zhai2022scaling, ; wen2024understanding, ; hu2024minicpm, ).
Across all three LRSs, we provide a detailed analysis of how the scaling exponents depend on task difficulty and model capacity. Our results show that compute-optimal allocation consistently favors data over model parameters, with the precise trade-off determined by the interplay between task difficulty and model capacity. Crucially, higher-capacity models are uniformly more compute-efficient across all settings considered. These findings offer a theoretical explanation for the widely adopted practice in scaling LLM training hoffmann2022training .
Finally, we also demonstrate the practical relevance of our FSL by applying it to LLM pre-training scenarios, providing theoretical insights into the empirical success of multiple-power-law heuristics proposed in luo2024a .
From a technical perspective, our FSL is derived by modeling online SGD using an intrinsic-time stochastic differential equation (SDE). A key innovation in our analysis is the introduction of the concept of intrinsic time, which serves as a fundamental quantity for capturing the scaling behavior induced by learning-rate schedules. We believe that both the intrinsic-time SDE framework and the resulting FSL provide a versatile foundation for future investigations into the theoretical principles underlying learning rate scheduling and LLM pre-training dynamics.
@@ 1.2 Related Work
Among the growing body of work seeks to theoretically explain scaling laws (sharma2020neural, ; hutter2021learning, ; maloney2022solvable, ; wei2022more, ; jain2024scaling, ; michaud2024quantization, ; nam2024exactly, ; atanasov2024scaling, ; dohmatob2024tale, ; bahri2024explaining, ; bordelon2024dynamical, ; lin2024scaling, ; paquette2024fourplus, ; bordelon2024feature, ), the most closely related to our work are (bordelon2024dynamical, ; paquette2024fourplus, ; lin2024scaling, ; bordelon2024feature, ), which also analyze scaling laws for teacher–student kernel regression. (bordelon2024dynamical, ) consider full-batch gradient flow, while (paquette2024fourplus, ) and (bordelon2024feature, ) focus on online SGD. A key limitation of these studies is their omission of label noise, which models irreducible risk and plays a critical role in determining compute-optimal scaling. (lin2024scaling, ) considers exponential-decay LRS (ge2019step, ), but do not account for the impact of batch size tuning. Building on this line of work, we propose a unified FSL that applies to general LRSs, which only recovers prior results as special cases but also substantially extends them by capturing the compounding effects of LRS and label noise.
LRS-aware scaling laws.
Recently, tissue2024scaling heuristically proposed a momentum law to predict loss trajectories in LLM pre-training. Building on this, luo2024a introduced a heuristic multi-power law that demonstrates improved empirical performance. These empirical studies highlight the existence of scaling laws that explicitly depend on LRS. In this work, we provide the first theoretical explanation for this phenomenon, establishing a principled foundation for LRS-aware scaling behavior.
2 Preliminaries
@@
2.1 Teacher–Student Kernel Regression
In this paper, we build upon the teacher–student kernel regression, a widely adopted, analytically tractable setup for studying scaling laws (maloney2022solvable, ; bahri2024explaining, ; bordelon2024dynamical, ; paquette2024fourplus, ; lin2024scaling, ; bordelon2024feature, ). In contrast to previous works, we provide an explicit analysis of how the LRS influences the scaling behavior. Formally, let $\mathbf{x}\in\mathbb{R}^{d}$ denote the data input that follows a distribution $\mathcal{D}$. The observed label is generated by $y:=f^{*}(\mathbf{x})+\epsilon,$ where $f^{*}:\mathbb{R}^{d}\to\mathbb{R}$ is the target function and $\epsilon$ denotes the label noise. We assume that $\epsilon\sim\mathcal{N}(0,\sigma^{2})$ and is independent of $\mathbf{x}$. The target function is assumed to be given by a teacher model:
$f^{*}(\mathbf{x}):=\langle\bm{\bm{\phi}}(\mathbf{x}),\bm{\theta}^{*}\rangle,$
where the target weight $\bm{\theta}^{*}\in\mathbb{R}^{N}$ with $N\in\mathbb{N}_{+}\cup\{\infty\}$, and $\bm{\bm{\phi}}(\cdot)=(\phi_{1}(\cdot),\phi_{2}(\cdot),\cdots,\phi_{N}(\cdot))^{\top}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{N}$ represents a feature map. We also assume the covariance matrix
$\mathbf{H}:=\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}[\bm{\bm{\phi}}(\mathbf{x})\bm{\bm{\phi}}(\mathbf{x})^{\top}]=\operatorname{diag}\{\lambda_{1},\lambda_{2},\cdots,\lambda_{N}\}$
with eigenvalues $\lambda_{1}\geqslant\lambda_{2}\cdots\geqslant\lambda_{N}$. This diagonal assumption is made purely for simplicity of presentation, and all our results can be straightforwardly extended to the non-diagonal case.
Assumption 2.1 (Hypercontractivity condition).
@@ 2.1 Teacher–Student Kernel Regression
There exists constants $C_{1},C_{2}>0$ such that it holds for any p.s.d. matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ that
$C_{1}\mathrm{tr}(\mathbf{H}\mathbf{A})\mathbf{A}\preceq\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}[\bm{\bm{\phi}}(\mathbf{x})^{\top}\mathbf{A}\bm{\bm{\phi}}(\mathbf{x})\cdot\bm{\bm{\phi}}(\mathbf{x})\bm{\bm{\phi}}(\mathbf{x})^{\top}]-\mathbf{H}\mathbf{A}\mathbf{H}\preceq C_{2}\mathrm{tr}(\mathbf{H}\mathbf{A})\mathbf{A}.$
@@
This hypercontractivity condition mei2022generalization essentially requires that the fourth-order moments of $\bm{\bm{\phi}}(\mathbf{x})$ can be controlled by its second-order moments. It is satisfied, for example, by Gaussian features $\bm{\phi}(\mathbf{x})\sim\mathcal{N}(0,\mathbf{H})$, as justified by Lemma E.1.
To learn $f^{*}$, we consider a finite-width student model defined by
$f(\mathbf{x};\mathbf{v}):=\langle\mathbf{v},\mathbf{W}\bm{\bm{\phi}}(\mathbf{x})\rangle,$ (1)
@@
where $\mathbf{W}\in\mathbb{R}^{M\times N}$ represents the projection matrix that projects the $N$-dimensional features into a $M$-dimensional subspace and $\mathbf{v}\in\mathbb{R}^{M}$ denotes the trainable weights. Here $M$ can be interpreted as the model size. Concretely, we consider two choices of projector $\mathbf{W}$:
  • Top-$M$ Features. $\mathbf{W}=[\mathbf{I}_{M},\bm{0}_{M\times(N-M)}]$. In this case, the student model essentially selects the top-$M$ eigenfunctions $\{\phi_{j}\}_{j=1}^{M}$ as the projected features,
  • Random Features. Each entry of $\mathbf{W}$ is i.i.d. sampled from $\mathcal{N}(0,1/M)$.
Note that top-$M$ features gives a theoretically convenient setting that simplifies our analysis – widely adopted in prior studies of scaling laws to facilitate tractable analysis (nam2024exactly, ; ding2025scaling, ). For random features (bahri2024explaining, ; bordelon2024dynamical, ; paquette2024fourplus, ; lin2024scaling, ; bordelon2024feature, ), we will show that, in certain learning scenarios, their scaling behavior closely mirrors that of the top-$M$ case.
2.2 The Model Capacity and Task Difficulty
We now formalize our key notions, model capacity and task difficulty, that will be pivotal in our subsequent scaling-law analysis. Let $e_{j}:=\phi_{j}/\lambda_{j}^{1/2}$ for $j\in[N]$. Then, $\{e_{j}\}_{j=1}^{N}$ forms an orthonormal basis in $L^{2}(\mathcal{D})$. When $M=\infty$, the student’s hypothesis space coincides with the reproducing‑kernel Hilbert space (RKHS) $\mathcal{H}_{k}$ associated with the kernel $k(\mathbf{x},\mathbf{x}^{\prime}):=\langle\bm{\phi}(\mathbf{x}),\bm{\phi}(\mathbf{x}^{\prime})\rangle=\sum_{j=1}^{N}\lambda_{j}e_{j}(\mathbf{x})e_{j}(\mathbf{x}^{\prime})$ (bach2024learning, ) (see a detailed discussion in Section C.1.1.

### v3

# Functional Scaling Laws in Kernel Regression: Loss Dynamics and Learning Rate Schedules

@@ Abstract
Scaling laws have emerged as a unifying lens for understanding and guiding the training of large language models (LLMs). However, existing studies predominantly focus on the final-step loss, leaving open whether the entire loss dynamics obey similar laws and, crucially, how the learning rate schedule (LRS) shapes them. We address these gaps in a controlled theoretical setting by analyzing stochastic gradient descent (SGD) on a power-law kernel regression model. The key insight is a novel intrinsic-time viewpoint, which captures the training progress more faithfully than iteration count. We then establish a Functional Scaling Law (FSL) that captures the full loss trajectory under arbitrary LRSs, with the schedule’s influence entering through a simple convolutional functional. We further instantiate the theory for three representative LRSs—constant, exponential decay, and warmup–stable–decay (WSD)—and derive explicit scaling relations in both data- and compute-limited regimes. These comparisons explain key empirical phenomena: (i) higher-capacity models are more data- and compute-efficient; (ii) learning-rate decay improves training efficiency; and (iii) WSD-type schedules outperform pure decay. Finally, experiments on LLMs ranging from 0.1B to 1B parameters demonstrate the practical relevance of FSL as a surrogate model for fitting and predicting loss trajectories in large-scale pre-training.
@@ 1 Introduction
It is well established that the training of large-scale deep learning models mysteriously follows scaling laws, whereby model performance improves predictably with available resources such as compute or data [20]. In particular, the landmark study by Kaplan et al. [27] demonstrated that, in LLM pre-training, the loss $L$ decreases with model size $M$ and dataset size $D$ according to a remarkably simple power-law relation:
$L(M,D)=L_{0}+C_{M}M^{-\alpha_{M}}+C_{D}D^{-\alpha_{D}},$ (1)
where $\alpha_{M}$ and $\alpha_{D}$ are the scaling exponents, $L_{0}$ denotes the irreducible loss, and $C_{M},C_{D}$ are some constants. Such empirical relations have proven robust across model architectures, scales, and training setups [21, 61, 38], and have become foundational principles for guiding LLM development [19, 26, 1, 5, 56, 29]. In practice, they are now routinely used to design optimal resource-allocation strategies [21] and to tune key hyperparameters such as learning rates and batch sizes [38, 31].
Despite their empirical success, the theoretical understanding of scaling laws remains limited. Recent studies have begun to illuminate the underlying mechanisms [55, 23, 41, 64, 25, 44, 45, 2, 15, 3, 8, 37, 51, 9, 72], yet two important gaps persist:
  • Determinants of scaling efficiency. Existing studies lack a systematic characterization of how key factors—such as model capacity, task difficulty, and hyperparameter choices—govern scaling efficiency, as reflected by the exponents $\alpha_{M}$ and $\alpha_{D}$. In particular, learning rate schedules (LRSs) are known to be critical in practice [42, 4, 18], but their precise role in shaping scaling efficiency remains unclear.
  • Beyond the final-step loss. The classical scaling law (1) addresses only the end-of-training loss [27, 21], leaving open whether the entire loss trajectory exhibits analogous scaling behavior. While recent empirical evidence [60, 40] suggests such possibility, a systematic study and theoretical justification are still missing.
Figure 1: Functional Scaling Law (FSL) accurately captures the loss dynamics and scaling behavior of SGD in kernel regression. In both subplots, solid lines denote the results of SGD, and dashed lines indicate the corresponding FSL predictions. (a) Loss dynamics of SGD (averaged over 1000 runs) compared with FSL predictions under three learning-rate schedules: cosine, WSD-like, and a non-standard cyclic schedule. (b) Final-loss scaling predicted by FSL using the analytical formulas from Section 5, compared with the mean of 200 independent SGD runs.
1.1 Our Contribution
In this paper, we take a step toward addressing these gaps in a controlled yet representative theoretical setting. We study stochastic gradient descent (SGD) training of the power-law kernel (PLK) regression—a widely adopted surrogate for scaling-law analysis [8, 3, 51, 37, 9]. The PLK regression is characterized by four parameters: the task difficulty $s$, the capacity exponent $\beta$, the model size $M$, and the label-noise level $\sigma$. To capture the influence of learning-rate schedules (LRSs), we model SGD via an intrinsic-time SDE, where the concept of intrinsic time emerges as a key quantity enabling a unified characterization of how different LRSs shape the loss dynamics and scaling behavior. Building on this formulation, we establish the Functional Scaling Law (FSL), which provides an accurate characterization of the entire loss dynamics beyond the traditional final-loss prediction.
Concretely, for a general intrinsic-time LRS $\gamma:[0,\infty)\!\to\![0,\infty)$, and under some conditions, the dynamics of the expected loss $\mathbb{E}[\mathcal{R}(\bm{\nu}_{t})]$ (where $t$ denotes the intrinsic time) satisfies:
$\mathbb{E}[\mathcal{R}(\bm{\nu}_{t})]-\underbrace{\frac{\sigma^{2}}{2}}_{\text{irreducible error}}\;\eqsim\;\underbrace{\frac{1}{M^{s\beta}}}_{\text{approx. error}}+\underbrace{e(t)}_{\text{signal learning}}+\underbrace{\!\int_{0}^{t}\!\mathcal{K}(t-z)\,[e(z)+\sigma^{2}]\,\gamma(z)\,\mathop{}\!\mathrm{d}z}_{\text{noise accumulation}},$ (2)
where $e(t)=(1+t)^{-s}$ and $\mathcal{K}(t)=(1+t)^{-(2-1/\beta)}$. Each term in FSL admits clear interpretation: $\tfrac{\sigma^{2}}{2}$ denotes the irreducible error caused by label noise, $M^{-s\beta}$ represents the approximation error, $e(t)$ characterizes the signal-learning dynamics under noiseless (full-batch) gradient descent, and the final term captures the injection and dissipation of gradient noise, with the LRS $\gamma$ entering through a tractable convolutional functional. The function $\mathcal{K}$, referred to as forgetting kernel, quantifies how fast the injected noise dissipates during training. Figure 1(left) shows that the FSL accurately captures the loss dynamics of SGD under diverse LRSs.
Building on FSL, we further derive explicit scaling relations for the final-step loss under three representative LRSs—constant, exponential decay [16], and warmup–stable–decay (WSD) [71, 22]—in both data-limited and compute-limited regimes. The results, summarized in Table 1 and Figure 1(right), recover and extend prior analyses [8, 9, 51, 37], while revealing several unifying insights.
  • Scaling efficiency of different schedules. WSD achieves the best scaling efficiency, followed by exponential decay and then constant schedules. This efficiency hierarchy provides theoretical justification for the importance of learning-rate decay and explains empirical success of WSD [71, 22, 59, 38].
  • Role of model capacity. Higher-capacity models are consistently more efficient in both compute and data, highlighting the necessity of scaling model capacity [27].
  • Data–model trade-off. Compute-optimal training requires scaling data more than model size, consistent with established heuristics in LLM pre-training [21].
  • Scaling law for peak learning rate. Optimal scaling requires the peak learning rate (LR) to scale appropriately with the training budget (data or compute), revealing the importance of careful peak LR tuning [7, 31].
Beyond PLK regression, we also apply the FSL ansatz to fit and predict loss trajectories from LLM pre-training experiments with model sizes ranging from 0.1B to 1B parameters, covering both dense and MoE architectures. These results highlight the potential of FSL as a practical surrogate for understanding and guiding LLM pre-training.
@@ 1.1 Our Contribution
Table 1: Learning-rate schedule (LRS) strongly influences scaling efficiency in power-law kernel regression. Efficiency is determined by two key factors: relative task difficulty $s\in(0,\infty)$ and model capacity $\beta>1$. We distinguish between an easy-learning regime ($s\geqslant 1-1/\beta$) and a hard-learning regime ($s<1-1/\beta$).
@@
Exponential-decay $D^{-\frac{s\beta}{1+s\beta}}(\log D)^{\frac{s\beta}{1+s\beta}}$ $D^{-s}(\log D)^{s}$ $C^{-\frac{s\beta}{2+s\beta}}(\log C)^{\frac{s\beta}{2+s\beta}}$ $C^{-\frac{s\beta}{1+{\beta}}}(\log C)^{\frac{s\beta}{1+{\beta}}}$
@@ 1.2 Related Work
Among the growing body of work seeking to theoretically explain scaling laws [55, 23, 41, 64, 25, 44, 45, 2, 15, 3, 8, 37, 51, 9], the most closely related are [8, 51, 9, 37], which also analyze PLK regression (often written in the equivalent linear-regression form). Specifically, [8] studies gradient flow, [51, 9] analyze SGD with a constant LRS, and [37] considers an exponential-decay LRS. In contrast, we establish a unified scaling law applicable to general LRSs, which not only recovers these prior results as special cases but also substantially extends them by capturing the loss dynamics rather than only the final-step loss. This unification is enabled by introducing the key notion of intrinsic time, which more faithfully captures the effective training progress than the raw number of training steps.
Predicting loss trajectories in LLM pre-training.
Recent empirical studies [60, 40] have shown that the entire loss trajectory of LLM pre-training—not merely the final loss—can be accurately captured by suitable scaling relations. A detailed description of the corresponding fitting procedures is provided in Appendix A.1. Our theory offers a theoretical explanation for these empirical findings. Interestingly, the multi-power-law (MPL) model proposed by [40] is closely connected to our FSL: through an integration-by-parts transformation, the FSL expression can be recast into a form that is nearly equivalent to the MPL formulation (see Appendix B.2).
Warmup-Stable-Decay (WSD) LRS.
A WSD schedule [71, 22] maintains a constant learning rate for a long stable phase, followed by a learning rate decay only near the end of training. Although unconventional, WSD has become popular in LLM pre-training [22, 18] and is already deployed in training industry-scale LLMs such as DeepSeek-V3 [38] and Kimi-K2 [59]. Yet its mechanism remains poorly understood. While recent works [65, 54] offer partial insights, we show—perhaps surprisingly—that even quadratic optimization, corresponding to a kernel regression problem, already reproduces the essential advantage of WSD. Furthermore, we quantify this advantage through explicit comparisons of scaling efficiency against constant and exponential-decay schedules.
1.3 Notation
@@
2 Power-Law Kernel (PLK) Regression
Let $\mathcal{X}$ denote the input domain and $\mathcal{D}$ the input distribution, and assume labels are generated by
$y=\langle\bm{\phi}(\mathbf{x}),\bm{\theta}^{*}\rangle+\epsilon,$
where $f^{*}(\mathbf{x}):=\langle\bm{\phi}(\mathbf{x}),\bm{\theta}^{*}\rangle$ is the target function, and the label noise $\epsilon\sim\mathcal{N}(0,\sigma^{2})$ is independent of $\mathbf{x}$. Here $\bm{\phi}:\mathcal{X}\to\mathbb{R}^{N}$ with $N\in\mathbb{N}_{+}\cup\{\infty\}$ is a feature map, satisfying the following assumption:
Assumption 2.1 (Hypercontractivity).
@@ 2 Power-Law Kernel (PLK) Regression
Let $\mathbf{H}:=\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}[\bm{\phi}(\mathbf{x})\bm{\phi}(\mathbf{x})^{\top}]$ be the feature covariance. There exist $\bm{\phi}$-dependent constants $\rho_{-},\rho_{+}>0$ such that for any PSD matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$,
$\rho_{-}\,\mathrm{tr}(\mathbf{H}\mathbf{A})\,\mathbf{A}\;\preceq\;\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}\!\Big[\left(\bm{\phi}(\mathbf{x})^{\top}\mathbf{A}\bm{\phi}(\mathbf{x})\right)\,\bm{\phi}(\mathbf{x})\bm{\phi}(\mathbf{x})^{\top}\Big]-\mathbf{H}\mathbf{A}\mathbf{H}\;\preceq\;\rho_{+}\,\mathrm{tr}(\mathbf{H}\mathbf{A})\,\mathbf{A}.$
@@
This condition ensures that the feature distribution is sufficiently regular—its fourth-order moments are controlled by the second-order ones [43]. It holds, for example, for Gaussian features $\bm{\phi}(\mathbf{x})\sim\mathcal{N}(0,\mathbf{H})$ with $\rho_{-}=1,\rho_{+}=2$ (see Lemma E.1).
Assumption 2.2.
$\mathbf{H}=\operatorname{diag}(\lambda_{1},\lambda_{2},\dots,\lambda_{N})$ with $\lambda_{1}\geqslant\lambda_{2}\geqslant\cdots\geqslant\lambda_{N}$.
This diagonalization assumption is made for simplicity. Since SGD is invariant under orthogonal rotations of the feature space, the analysis under this assumption is equivalent to the general case up to rotation.
@@
To learn $f^{*}$, we consider a model of width $M$: $f(\mathbf{x};\mathbf{v})=\sum_{j=1}^{M}v_{j}\mathbf{w}_{j}^{\top}\bm{\phi}(\mathbf{x})=:\langle\mathbf{v},\mathbf{W}\bm{\phi}(\mathbf{x})\rangle,$ where $\mathbf{v}\in\mathbb{R}^{M}$ denotes trainable weights and $\mathbf{W}\in\mathbb{R}^{M\times N}$ projects the $N$-dimensional features onto an $M$-dimensional subspace. We study two choices of projection $\mathbf{W}$:
  • Top-$M$ features: $\mathbf{w}_{j}=\mathrm{\mathbf{e}}_{j}$ for $j\in[M]$, i.e., selecting the top-$M$ features $\{\phi_{j}\}_{j=1}^{M}$;
  • Random-$M$ features: $\mathbf{w}_{j}\sim\mathcal{N}(0,I_{N})$ independently for $j\in[M]$.
The top-$M$ setting is a particularly simple yet analytically representative case, also adopted in prior scaling-law studies [45, 14]. For random features [3, 8, 51, 37, 9], we will show that, in certain regimes, their scaling behavior closely parallels that of the top-$M$ case. As clarified in Appendix A.3, our setup is equivalent to learning with the kernel $K_{\phi}(\mathbf{x},\mathbf{x}^{\prime}):=\bm{\phi}(\mathbf{x})^{\top}\bm{\phi}(\mathbf{x}^{\prime})$.
2.1 Model Capacity and Task Difficulty
We now formalize the key notions of model capacity and task difficulty. Let $\widehat{\phi}_{j}:=\phi_{j}/\lambda_{j}^{1/2}$ for $j\in[N]$, so that $\{\widehat{\phi}_{j}\}_{j=1}^{N}$ forms an orthonormal basis of $L^{2}(\mathcal{D})$.